# Intelligence at the Edge of Chaos

### v2

@@ 6 Related Work
Elementary Cellular Automata (ECA) serve as foundational models for exploring complexity arising from simple rules. Early research demonstrated that even minimalistic CA systems, governed by basic and local interaction rules, could generate intricate patterns over time (Wolfram, 1983). Wolfram’s extensive investigations into one-dimensional CA revealed that certain rules produce behaviors ranging from stable and periodic to chaotic and complex (Wolfram, 1984). In his work Cellular Automata and Complexity, Wolfram classified one-dimensional CA into four classes based on their dynamic behaviors, highlighting the rich complexity that simple systems can exhibit (Wolfram & Mallinckrodt, 1994).
@@ 7 Discussion
Our findings demonstrate that large models are capable of learning complex, non-trivial solutions even when simpler, more trivial ones are available, likely because they are overparametrized. For lower-complexity ECA rules, the models often adopt a trivial solution, focusing only on the current state since no history is needed. However, when exposed to more complex rules, models tend to leverage prior states, as indicated by the attention patterns in Figure 4. Despite the memoryless nature of ECA systems, overparameterized models explore a broader search space, and solutions that integrate historical information may be more robust. We hypothesize that by learning to incorporate past states, the model develops generalizable logic that can be reused across tasks. In contrast, a model relying solely on a trivial, state-specific rule would struggle to transfer its knowledge to more complex downstream tasks. Thus, the ability to learn from past states may be key to the model’s success in adapting to diverse problems.
@@
Appendix A Visualizing Learned Representations Using CKA Similarities
Figure 6: UMAP projection of CKA similarities between models trained on ECA rules. Each point represents a model, colored by its corresponding rule complexity (Lempel-Ziv, Compression, Lyapunov, and Krylov complexities). Models trained on rules with similar complexities cluster together, indicating that they learn similar representations. Chaotic rules like Rule 105 and 150 are notably closer to models trained on lower complexity rules, while models trained on complex, non-chaotic rules cluster more distinctly.
Figure 7: UMAP projection of CKA similarities between models trained on ECA rules, annotated with the rule number.
To investigate the learned representations of the pre-trained models, we use center kernel alignment (CKA) to compute similarities between each pair of models. CKA is a method for comparing representations learned by neural networks, measuring how well the internal weight matrices of two models align (Kornblith et al., 2019). In our analysis, we compute pairwise similarities across all models trained on different ECA rules, with complexity levels ranging from low to high.
To visualize these similarities, we project the CKA similarity embeddings into two dimensions using UMAP, a popular dimensionality reduction technique. The resulting plots, shown in Figures 6 and 7, reveal that models trained on rules with similar complexity tend to cluster together, indicating that they have learned similar internal representations.
Interestingly, models trained on highly chaotic rules, such as Rule 105 and Rule 150, are located closer to models trained on lower-complexity rules, suggesting that despite the high complexity of their training data, their learned representations are more akin to those trained on simpler dynamics. In contrast, models trained on complex but not chaotic rules tend to form distinct clusters, indicating that they develop more unique internal structures compared to models trained on either simple or highly chaotic rules.
These results suggest that complexity plays a significant role in shaping model representations. Models trained on rules with moderate complexity tend to develop distinct, well-formed internal representations, while chaotic rules may push models toward more generic representations, aligning them with simpler rules.
Appendix B Long vs Short Term Prediction with other complexity measures
Figure 8: Efficiency comparison between models trained on short-term and long-term prediction tasks, colored by the other complexity measures. We observe the same trend as with Lempel-Ziv complexity in Figure 5.
Figure 8 presents an efficiency comparison between models trained on short-term (1-step) and long-term (5-step) prediction tasks akin to Figure 5, but with the remaining complexity measures. We observe the same trend as with the Lempel-Ziv complexity. Models tend to perform better when trained on shorter prediction horizons, challenging the assumption that longer prediction horizons lead to better understanding of underlying structures.

### v3

@@ 6 Related Work
Elementary Cellular Automata (ECA) serve as foundational models for exploring complexity arising from simple rules. Early research demonstrated that even minimalistic CA systems, governed by basic and local interaction rules, could generate intricate patterns over time (Wolfram, 1983). Wolfram’s extensive investigations into one-dimensional CA revealed that certain rules produce behaviors ranging from stable and periodic to chaotic and complex (Wolfram, 1984). In his work Cellular Automata and Complexity, Wolfram classified one-dimensional CA into four classes based on their dynamic behaviors, highlighting the rich complexity that simple systems can exhibit (Wolfram & Mallinckrodt, 1994). Wolfram’s Principle of Computational Equivalence posits that systems with sufficiently complex behavior can exhibit computational capabilities equivalent to universal computation (Wolfram & Gad-el Hak, 2003). ECA Rule 110, which has been proven to be Turing complete (Cook et al., 2004), demonstrates that even simple, rule-based systems can perform any computation given the right initial conditions. Recent works have explored the use of transformers to predict cellular automata states (Berkovich & Buehler, 2024).
@@
Reservoir Computing and Chaotic Systems
Reservoir Computing (RC) leverages fixed, high-dimensional dynamical systems—often randomly initialized recurrent neural networks—to model nonlinear and chaotic systems (Maass et al., 2002). It projects input signals into a high-dimensional space, capturing temporal patterns through inherent dynamics while only training the output weights (Lukoševičius & Jaeger, 2009). This simplifies training and allows the system to handle chaotic behavior without adjusting internal weights. RC has been successfully applied to tasks like time series prediction and system identification (Pathak et al., 2018), effectively harnessing chaotic system dynamics for computation.
@@ 7 Discussion
Our findings demonstrate that large models are capable of learning complex, non-trivial solutions even when simpler, more trivial ones are available, likely because they are overparametrized. For ECA rules generating lower-complexity data, the models often adopt a trivial solution, focusing only on the current state since no history is needed. However, when exposed to more complex data, models tend to leverage prior states, as indicated by the attention patterns in Figure 3. Despite the memoryless nature of ECA systems, overparameterized models explore a broader search space, and solutions that integrate historical information may be more robust. We hypothesize that by learning to incorporate past states, the model develops generalizable logic that can be reused across tasks. In contrast, a model relying solely on a trivial, state-specific rule would struggle to transfer its knowledge to more complex downstream tasks. Thus, the ability to learn from past states may be key to the model’s success in adapting to diverse problems.
@@
8 Reproducibility Statement
To facilitate the reproduction of our results, we have provided detailed descriptions of our data processing methods and experimental procedures in Sections 3 and 4. We utilized the GPT-2 model obtained from Hugging Face for our experiments. The complete code for our paper is now available in the following GitHub repository: https://github.com/vandijklab/Intelligence_at_the_edge_of_chaos.
Appendix A Scaling Experiments
In this section, we explore the interplay between data complexity, model capacity, and data volume in the learning process. We conducted scaling experiments using three transformer-based models:
  1. 1.
    A 67k-parameter custom single-layer transformer with one attention head and 64-dimensional embeddings (“Tiny”),
  2. 2.
    An 85M-parameter GPT-2 small model with 12 layers, 12 attention heads, and 768-dimensional embeddings (“Small”), and
  3. 3.
    A 708M-parameter GPT-2 large model with 36 layers, 20 attention heads, and 1280-dimensional embeddings (“Large”).
Each model is pretrained on various quantities of ECA-generated data with varying complexities.
Figure 5: Scaling experiments with varying quantities of data and different model sizes. Left: Number of tokens seen before convergence during pre-training for the Tiny model. Center: Number of tokens seen before convergence during pretraining for the Small model. Right: Validation loss as a function of token consumption for models trained on data from ECA Rule 110. Larger models achieve lower validation loss with fewer tokens, highlighting the improved data efficiency for larger models.
Figure 5 illustrates the results of these experiments. The left and middle panels show that as the complexity of the data increases, both models require more tokens to reach convergence. This trend indicates that higher data complexity demands more data for effective learning. Notably, the right panel demonstrates that the Small model achieves lower validation loss faster than the Tiny model when trained on complex data (and similarly for Large vs. Small), suggesting that larger models can learn more efficiently from complex patterns.
These findings highlight a trade-off between data complexity, model capacity, and data volume. While complex training data can enhance model capabilities, it necessitates more data for smaller models to learn effectively. However, increasing model capacity can mitigate this requirement by enabling faster convergence, thereby reducing the amount of data needed.
Appendix B The Role of Temporal Structure in Learning Representations
To discern whether temporal structure is necessary or if spatial complexity alone is sufficient for better feature learning, we performed an ablation study by disrupting the sequential nature of the ECA data. Specifically, we randomly shuffled the temporal order of the states generated by the ECAs while preserving their spatial configurations. This manipulation retains the spatial complexity but removes the inherent temporal structure, effectively isolating the spatial component.
We trained models on this temporally shuffled data following the same procedures as detailed in Section 3. The performance of these models on the easy reasoning task showed a significant drop compared to models trained on the original, temporally ordered data. As shown in Figure 6, models trained on temporally structured data reached 80% accuracy in significantly fewer epochs than those trained on temporally shuffled data.
These results underscore the critical role of temporal information in ECA data in enabling models to learn representations that generalize effectively. While spatial complexity provides rich patterns, it is the structured temporal evolution that allows models to capture dynamic relationships and develop more sophisticated reasoning abilities. This experiment confirms that both spatial and temporal complexities are necessary components for training models capable of generalizing and performing complex reasoning tasks.
Figure 6: Impact of temporal shuffling on downstream task performance. Models trained on temporally shuffled ECA data exhibit decreased efficiency on the easy reasoning task compared to models trained on ordered data, highlighting the importance of temporal structure in learning effective representations. The average efficiency for each Wolfram class on ordered data is shown with dashed lines for reference. ()(*)( ∗ ) indicates models that did not reach the 80% accuracy threshold.
Appendix C Downstream Task: Nim Game
To further evaluate the models’ ability to generalize and perform strategic reasoning, we tested their performance on predicting optimal moves in the game of Nim, a classic mathematical strategy game (Bouton, 1901). In Nim, players take turns removing objects from distinct heaps, aiming to be the one who removes the last object. Optimal play involves computing the binary addition of heap sizes, making this game an ideal test for logical reasoning and pattern recognition.
We constructed a dataset comprising sequences of Nim game states and their corresponding optimal moves. To produce diverse scenarios for a challenging task, we simulate games starting with between 5 and 15 heaps and 1 to 10 items in each heap. Each game state is represented as a string starting with the initial configuration of the heaps, followed by a sequence of moves until the game terminates when all the heaps are empty. Each move is encoded as the player number, followed by the heap identifier, followed by the number of items removed. For example, “1B3” signifies “Player 1 removed 3 items from heap B”. Each possible move and heap state corresponds to a unique token for the embedding layer of the model. The dataset was split into training and validation sets using a 90-10 ratio.
Using the same approach as in previous downstream tasks, we trained models with the ECA-pretrained weights frozen, allowing only the input embedding and output layers to be updated. Figure 7 shows the relationship between model efficiency and the complexity of the ECA rule used during pretraining. We found that the Nim game is considerably harder than the ARC reasoning tasks, and most models are unable to get accuracies greater than 20%. We adjust our definition of “efficiency” accordingly to 1 divided by number of epochs to reach 20% accuracy. We observed a significant ($p<0.05$) positive correlation between data complexity and model performance on the Nim task, consistent with our earlier findings on reasoning and chess benchmarks. Models pretrained on more complex data required fewer epochs to reach a given level of accuracy, indicating that complex pretraining data facilitates the acquisition of strategic reasoning abilities.
These results reinforce the idea that pretraining on complex data structures enables models to develop more sophisticated representations, which can be effectively leveraged in downstream tasks.
Figure 7: The relationship between performance in the Nim game task and data complexity. Higher data complexity is correlated with increased efficiency, measured as 1 divided by the time to reach 20% accuracy. The data has a statistically significant correlation coefficient of $r=0.63$, consistent with the trends we observed with the other downstream tasks.
Appendix D Correlation Coefficients and Hypothesis Testing
We computed the Pearson correlation and the distance correlation coefficients (Edelmann et al., 2021) between the rule complexity and the performance on all downstream tasks. Whereas the Pearson correlation coefficient measures the linear dependence between variables, the distance correlation is capable of measuring non-linear association as well. The correlation coefficients and their respective p-values for the Lempel-Ziv complexity are given in Table 1.
Table 1: Summary of Correlation Analysis Results for Downstream Tasks
Task Pearson Correlation Distance Correlation
ARC-Easy 0.7322 ($p=4.93\times 10^{-18}$) 0.7163 ($p=4.97\times 10^{-3}$)
ARC-Hard 0.3896 ($p=5.04\times 10^{-4}$) 0.4444 ($p=4.97\times 10^{-3}$)
Chess 0.4371 ($p=9.73\times 10^{-3}$) 0.4329 ($p=2.99\times 10^{-2}$)
NIM 0.6314 ($p=3.31\times 10^{-6}$) 0.7108 ($p=4.97\times 10^{-3}$)
Appendix E Results for other complexity measures
The downstream results for the ARC Easy, ARC Hard, and Chess tasks as a function of the compression complexity, Lyapunov exponent, and Krylov complexity measures are shown in Figure 8. We observe the same general patterns that we see with the Lempel-Ziv complexity in Figure 2.
Figure 8: Relationship between downstream task performance and data complexity for other complexity measures. Rows depict easy reasoning, hard reasoning, and chess move prediction tasks, while columns show compression complexity, Lyapunov exponent, and Krylov complexity, respectively. Baseline results for randomly initialized transformers are shown with a dashed black line.